a man is three times as old as his son. four years ago, he was four times as old as his son was at that time.how old is the son?
i think its twelve. i've done it in a trial and error way, but i need a solution..thank you^^

Found 2 solutions by solver91311, dannia22:
Answer by solver91311(24713)



Let represent the son's age now. Then the father's age now is . Four years ago the father's age was and that is four times what the son's age was four years ago, namely , so:



Solve for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by dannia22(1)

4(x-4)=3x-4
4x-16=3x-4
4x-3x=-4+16
x=12
thank you^^